A​ person's website specializes in the sale of rare or unusual vegetable seeds. He sells packets of​ sweet-pepper seeds for ​$2.16 each and packets of​ hot-pepper seeds for ​$4.40 each. He also offers a 16​-packet mixed pepper assortment combining packets of both types of seeds at ​$2.58 per packet. How many packets of each type of seed are in the​ assortment?
Answer by addingup(3677)

2.16x+4.40(16-x) = 2.58(16)
2.16x+70.40-4.40x = 41.28
-2.24x = -29.12
x 13
So, there are 13 packets of the sweet pepper seeds at 2.16 and 16-13 = 3 packets of the hot pepper seeds at 4.40